# Multi-SpaCE: Multi-Objective Subsequence-based Sparse Counterfactual Explanations for Multivariate Time Series Classification

### v1

@@ 1 Introduction
Machine Learning, and particularly Deep Learning, has become increasingly pervasive, automating tasks and improving decision-making across industries such as healthcare, finance, and manufacturing [1]. However, despite their benefits, Deep Learning systems often face critical challenges, such as their “black-box” nature. These systems excel in complex tasks but lack transparency, making it difficult for humans to interpret their decisions. This opacity hinders their deployment in high-stakes applications, where understanding the rationale behind decisions is crucial. To address this issue, eXplainable Artificial Intelligence (XAI) has emerged as a field dedicated to making machine learning models more interpretable, fair, and transparent [2]. By enabling decision-makers to trust AI systems, XAI aims to facilitate the integration of AI into critical domains.
@@
Among the many explanation techniques in XAI, counterfactual explanations (CFEs) [3] have gained significant attention due to their alignment with human cognitive processes [4]. CFEs identify the smallest changes required by an input instance to achieve a desired outcome, thus allowing users to explore “what-if” scenarios. Originally framed as optimization problems, CFEs searched for proximity between the original instance and the counterfactual, while ensuring the validity of the counterfactual (the change in the outcome) [5]. With time, counterfactual explanations have evolved to incorporate desiderata such as plausibility [6, 7], diversity [8], and actionability [9]. As new desiderata were identified, the conflicts between them also became apparent. Multi-objective Optimization has become an effective tool to address these trade-offs between conflicting properties [10, 11, 12], providing users with a Pareto front of solutions that can be easily adapted to their preferences.
@@
While CFEs were initially developed for tabular data, their adaptation to time series remains an active area of research. Existing methods typically leverage subsequences of changes to address the higher dimensionality and serial correlations inherent in time series inputs [13, 14, 15, 16, 17, 18, 19, 12, 20, 21]. However, most methods also enforce rigid constraints, such as single or fixed-length subsequences, limiting their flexibility and effectiveness. Furthermore, many approaches remain focused on univariate settings, failing to account for the complexities of the multivariate time series which are commonly found in real-world applications. Most importantly, validity is frequently treated as another objective rather than a strict requirement. While many works define validity as a constraint, either verbally or mathematically [5], the heuristic approaches used to solve these problems often fail to enforce it. We argue that validity is fundamental and should be treated as a strict requirement. Without validity, all the other performance metrics lose their significance, as they are no longer associated to explanations that achieve the desired outcome. Methods that fail to enforce validity might be unsuitable for practical applications.
@@
With this analysis, we argue that i) Multi-objective obtimization is desired, as it offers a natural framework for addressing conflicting properties; ii) methods should support multivariate time series data to ensure applicability to a wide range of real-world problems; and iii) the validity of counterfactual explanations should be treated as a strict requirement rather than just another metric to optimize.
@@
These principles form the foundation of our proposed method, Multi-SpaCE, which addresses these challenges by ensuring the validity of the returned counterfactuals, supporting multivariate time series, and leveraging a multi-objective optimization framework. Multi-SpaCE extends our previous work, Sub-SpaCE [21], that uses a genetic algorithm (GA) with custom mutation and initialization to generate counterfactual explanations by optimizing sparsity, plausibility and minimizing the number of subsequences. Multi-SpaCE inherits all the good properties of Sub-SpaCE (such as its model-agnostic nature, its computational efficiency or the validity of all the generated CFEs), while overcoming its two major limitations: its restriction to univariate data and its reliance on extensive experimentation to balance multiple desiderata. By introducing a multi-objective optimization framework based on Non-Dominated Sorting Genetic Algorithm II (NSGA-II) [22], Multi-SpaCE offers a Pareto-Front of solutions, enabling users to explore trade-offs between objectives. The optimization task identifies the points in the original input to modify, substituting their values with those of the Nearest Unlike Neighbor (NUN) [13]. Figure 1 presents the architecture of the proposed solution: the NUNs obtained for each input instance, the black box classifier to be explained, the NSGA algorithm used to generate candidate CFEs, the Autoencoder introduced to measure their plausibility, and the Pareto-front of counterfactuals returned by Multi-SpaCE.
Figure 1: Multi-SpaCE block diagram.
@@
To the best of our knowledge, Multi-SpaCE is the first multi-objective optimization method that ensures valid counterfactual explanations for multivariate time series classification problems. Extensive experimentation demonstrates its ability to achieve perfect validity across datasets while outperforming existing methods in multiple metrics.
@@ 2 Related work
The field of eXplainable Artificial Intelligence (XAI) has experienced significant growth in recent years, resulting in the development of numerous methods tailored to specific domains and data types [23]. However, XAI methods for time series have received considerably less attention than those designed for tabular or image data [24]. Many existing methods are adaptations of those designed for tabular or image data [25], and cannot properly cope with the unique characteristics of time-ordered inputs [26]. Time series data, such as electroencephalograms (EEG), electrocardiograms (ECG), or financial time series require domain expertise to interpret, making traditional methods less effective for generating meaningful explanations [24].
@@
This trend extends to counterfactual explanations. Formally, a counterfactual explanation $x^{\prime}$ is the smallest variation of an original instance $x$ that changes the predicted outcome $y=b(x)$ of a black-box classifier $b\!:X\rightarrow[0,1]$, subject to $b(x^{\prime})\neq y$ [5]. Most counterfactual explanation methods are designed for tabular data. When applied to time series, these methods may not scale well with the increased dimensionality of inputs and often fail to account for the serial correlations that characterize time series data [13].
@@
To address these challenges, methods such as (NG) [13] introduced contiguity, ensuring that changes occur as subsequences aligned with the sequential structure of time series. NG finds the Nearest Unlike Neighbot (NUN), the closest example to the instance being explained that belongs to a desired class, and substitutes the shortest subsequence of the original instance with the corresponding subsequence from the NUN to achieve the desired outcome. The main drawback of NG is that it is limited to univariate settings and allows only a single subsequence of changes, thus restricting its practical applicability. COMTE, by Ates et al. [14], extended this idea to multivariate time series by optimizing the number of channels substituted from the NUN to produce a valid counterfactual. However, this approach is still limited to a single subsequence per channel, which may introduce unnecessary changes and break the existing relationships between channels.
@@
Other methods leverage shapelets to guide counterfactual explanations. For example, Local Agnostic Subsequence-based Time Series Explainer (LASTS) [15, 16] uses the latent space of a Variational Autoencoder (VAE) to generate a set of exemplars and counterfactuals in the neighborhood of the instance to explain. It then derives classification rules based on the presence of specific shapelets. Another work, Shapelet Based Counterfactual Explanations (SETS) [27], extracts discriminative shapelets for each class. It then localizes them in the input to explain, and substitutes their corresponding values with those of the NUN.
@@
More recent approaches, such as Attention-Based Counterfactual Explanations (AB-CF) [18] and Discord-based Counterfactual Explanations (DiscoX) [19], address multivariate time series but rely on fixed-length subsequences. AB-CF greedily selects fixed-length multichannel subsequences that maximize classification entropy when isolated, while DiscoX identifies fixed-length discords (subsequences maximally distant from their nearest neighbors in the desired class) and replaces them with the nearest subsequences from the desired output class. Although these methods perform well in some scenarios, their reliance on fixed-length subsequences often results in unnecessary changes. Additionally, their heuristic and greedy approaches limit validity rates, reducing their applicability in real-world scenarios.
@@
Another line of research still follows the optimization-based counterfactual generation originally proposed by Wachter et al. [3]. Building on this foundational work, Glacier [17] generates counterfactuals in the latent space of an autoencoder while penalizing changes outside relevant subsequences identified by LIMESegment [28]. However, Glacier treats validity as a loss term rather than a strict constraint, leading to low validity rates, and remains limited to univariate datasets. Genetic algorithms (GAs) have also been applied to generate CFEs. TSEvo [12] introduced a multi-objective approach and diverse mutation operators to optimize both the location and magnitude of changes. However, TSEvo’s high-dimensional search space results in long execution times and low validity rates. More recently, TX-Gen [20] tackles multi-objective optimization by finding a single subsequence of changes. Instead of relying on the NUN, TX-Gen employs a compressed representation of the subsequence and uses a simple AR model to generate the content of the subsequences proposed by the optimizer. Although TX-Gen reports perfect validity, it was only tested in univariate settings and might struggle with plausibility in more complex scenarios, where a simple generative AR model could not provide enough flexibility. Finally, Sub-SpaCE [21] balances sparsity, plausibility, and the number of subsequences by framing counterfactual generation as a single-objective optimization problem solved with a GA with custom mutation and inicialization. Sub-SpaCE identifies a binary mask of changes, indicating the locations where the values of the original instance should be substituted with those of the NUN. While effective, Sub-SpaCE is restricted to univariate data and lacks a multi-objective framework, requiring experimentation to balance objective relevance.
@@
As stated above, contiguity is a desired property, as grouped changes in the form of subsequences enhance interpretability. However, current methods either promote contiguity indirectly (e.g., TSEvo by using the mutation operators), or limit their flexibility by imposing: i) a single subsequence of changes (e.g., NG, COMTE and TX-Gen) or; ii) fixed-length subsequences (e.g., AB-CF and DiscoX) that limit adaptability to complex scenarios. Additionally, methods should be prepared to work in multivariate settings, which are ubiquitous in real-world applications, something that NG, Glacier, TX-Gen, and Sub-SpaCE do not support. Finally, and most importantly, except for a few approaches (such as NG, TX-Gen and Sub-SpaCE), most methods fail to ensure perfect validity, treating it as another objective rather than as a strict constraint.
@@ 3.2.1 Extension to multivariate problems
In both cases the NUN is the same as in Sub-SpaCE: the instance, $\mathbf{x}^{nun}$, closest to $\mathbf{x}$ in Euclidean space. Both approaches have been tested, and the best results (see Appendix B) are obtained when both are used: the common mask across channels is used at the beggining of the optimization, and the independent mask is used to then improve the solution. The pseudocode of Multi-SpaCE is shown in Algorithm 1.
@@ 4.1 Setup
For each dataset in Table 1, we trained multiple black-box classifiers $b$ under different configurations to perform classification. Additionally, we trained several Autoencoders ($f_{AE}$), Isolation Forests ($f_{IF}$), and Local Outlier Factor ($f_{LOF}$) models to assess the plausibility of the generated counterfactuals. Details about the trained models are provided in Appendix A. We then selected the best-performing models of each family. The model achieving the highest weighted F1-score on the test set was selected as the classifier to be explained. For the Autoencoder, the one with the lowest reconstruction error was chosen. Models with the best silhouette scores were used for the Isolation Forest and Local Outlier Factor. Classifiers and autoencoders were implemented using the TensorFlow framework, while Isolation Forests and Local Outlier Factors were trained using the Scikit-Learn library.
The parameter values of Multi-SpaCE were empirically set and kept consistent across datasets to ensure robustness to multiple scenarios. Instead of extensive fine-tuning of all possible parameters, we focused on evaluating the influence of mutation parameters and the different types of masks proposed in Section 3.2.1, as they are central to our framework. Based on these evaluations (see Appendix B), the final Multi-SpaCE uses a population size $N=100$ and total number of generations $G=100$, with the first $G_{1}=75$ generations for optimizing a common mask across channels with $p^{e}=p^{c}=0.75$ and $p^{p}=0$, and the last $G_{2}=25$ generations for pruning unnecessary subsequences using the independent mask setting with $p^{p}=0.5$. We use an initialization percentage $h=20\%$, increased by $h_{inc}=20\%$ if no valid counterfactual is found after $G_{reinit}=50$ generations. The penalization term for incorrect classes was set to $\nu=100$ to strictly enforce valid counterfactuals. The implementation of Multi-SpaCE and experimental code is publicly available at https://github.com/MarioRefoyo/Multi-SpaCE.
@@
Unlike the baseline methods, which return a single counterfactual considered optimal, Multi-SpaCE generates a set of solutions that are optimal under different weightings of the objectives. To fairly compare Multi-SpaCE with the baselines, we use a custom utility function to assign weights to each objective: a weight of 0.1 to the adversarial objective ($o_{1}$), 0.3 to sparsity ($o_{2}$), 0.4 to the number of subsequences ($o_{3}$), and 0.2 to plausibility ($o_{4}$). These weights reflect our preferences and are aligned with the parameter values originally assigned to the objective function of Sub-SpaCE [21]. After applying the utility function, we select the counterfactual from the Multi-SpaCE solution set that maximizes the utility score. This selected counterfactual is then used to compute the metrics, with the results reported in Section 4.5.
@@ 4.2 Datasets
The performance of Multi-SpaCE was evaluated on datasets from the UCR archive [29] and the UEA archive [30]. For each dataset, the best classifier was selected based on its weighted F1-score on the test set. Only datasets where the classifier achieved an F1-score greater than 80% were considered, ensuring that counterfactuals were generated for reasonably accurate models.
Under the multivariate setting, only 10 datasets surpassed the 80% threshold. For the univariate setting, over 30 datasets met the criterion, out of which 15 were selected to ensure diversity based on the following criteria: i) number of classes; ii) time series length; and iii) different data characteristics. In this regard, half of the chosen datasets had a binary classification outcome, as some baseline methods are tailored for binary classification. Additionally, we selected extreme cases, such as NonInvasiveFatalECGThorax2 with a high number of classes. Datasets with varying lengths were included, such as ItalyPowerDemand (24 steps) and HandOutlines (2709 steps). Finally, we selected datasets with diverse characteristics. For instance, FordA was chosen due to frequencies being representative of output classes, while datasets like ECG200 and TwoPatterns were included as waveform patterns is what defines class separability. Table 1 provides a detailed listing of all datasets (train and test size, time series length, number of channels and classes), as well as the test F1-scores of the trained classifiers.
@@
TwoPatterns 1000 4000 128 1 4 1.000
Strawberry 613 370 235 1 2 0.981
Plane 105 105 144 1 7 0.981
ItalyPowerDemand (ItalyPower) 67 1029 24 1 2 0.960
CinCECGTorso 40 1380 1639 1 4 0.951
Gunpoint 50 150 150 1 2 0.946
Coffee 28 28 286 1 2 0.928
HandOutlines 1000 370 2709 1 2 0.920
FordA 3601 1320 500 1 2 0.918
NonInvasiveFatalECGThorax2 (NI-ECG2) 1800 1965 750 1 42 0.915
ECG5000 500 4500 140 1 5 0.899
ProximalPhalanxOutlineCorrect (PPOC) 600 291 80 1 2 0.899
CBF 30 900 128 1 3 0.897
ECG200 100 100 96 1 2 0.853
FacesUCR 200 2050 131 1 14 0.851
PenDigits 7494 3498 8 2 10 0.981
SelfRegulationSCP1 (SR-SCP1) 268 293 896 6 2 0.901
NATOPS 180 180 51 24 6 0.885
Epilepsy 137 138 206 3 4 0.878
PEMS-SF 267 173 144 963 7 0.869
ArticularyWordRecognition (AWR) 275 300 144 9 25 0.863
UWaveGestureLibrary (UWave) 120 320 315 3 8 0.852
Cricket 108 72 1197 6 12 0.843
RacketSports 151 152 30 6 4 0.841
BasicMotions 40 40 100 6 4 0.838

### v2

@@ 1 Introduction
Machine Learning in general, and particularly Deep Learning, is becoming increasingly pervasive, automating tasks and improving decision-making across industries such as healthcare, finance, and manufacturing [1]. However, despite their benefits, Deep Learning systems often face critical challenges, such as their “black-box” nature. These systems excel in complex tasks but lack transparency, making it difficult for humans to interpret their decisions. This opacity hinders their deployment in high-stakes applications, where understanding the rationale behind decisions is crucial. To address this issue, eXplainable Artificial Intelligence (XAI) has emerged as a field dedicated to making machine learning models more interpretable, fair, and transparent [2, 3]. By enabling decision-makers to trust AI systems, XAI aims to facilitate the integration of AI into critical domains.
@@
Among the many explanation techniques in XAI, counterfactual explanations (CFEs) [4] have gained significant attention due to their alignment with human cognitive processes [5, 6]. CFEs identify the smallest changes required by an input instance to achieve a desired outcome, thus allowing users to explore “what-if” scenarios. Originally framed as optimization problems, CFEs searched for proximity between the original instance and the counterfactual, while ensuring the validity of the counterfactual (the change in the outcome) [7, 8]. With time, counterfactual explanations have evolved to incorporate desiderata such as plausibility [9, 10, 11], diversity [12], and actionability [13]. As new desiderata were identified, the conflicts between them also became apparent. Multi-objective optimization has become an effective tool to address these trade-offs between conflicting properties [14, 15, 16], providing users with a Pareto front of solutions that can be easily adapted to their preferences.
@@
While CFEs were initially developed for tabular data, their adaptation to time series remains an active area of research. Existing methods typically leverage subsequences of changes to address the higher dimensionality and serial correlations inherent in time series inputs [17, 18, 19, 20, 21, 22, 23, 16, 24, 25]. However, most methods also enforce rigid constraints, such as single or fixed-length subsequences, limiting their flexibility and effectiveness. Furthermore, many approaches remain focused on univariate settings, failing to account for the complexities of the multivariate time series which are commonly found in real-world applications. Most importantly, validity is frequently treated as another objective rather than a strict requirement. While many works define validity as a constraint, either verbally or mathematically [7], the heuristic approaches used to solve these problems often fail to enforce it. We argue that validity is fundamental in CFEs and should be treated as a strict requirement. Without validity, all the other performance metrics lose their significance, as they are no longer associated with explanations that achieve the desired outcome. Methods that fail to enforce validity might be unsuitable for practical applications.
@@
With this analysis, we argue that i) Multi-objective obtimization is desired, as it offers a natural framework to address conflicting properties; ii) methods should support multivariate time series data to ensure applicability to a wide range of real-world problems; and iii) the validity of counterfactual explanations should be treated as a strict requirement rather than just another metric to optimize.
@@
These three principles form the foundation of our proposed method, Multi-SpaCE, which addresses these challenges by ensuring the validity of the returned counterfactuals, supporting multivariate time series, and leveraging a multi-objective optimization framework. Multi-SpaCE extends our previous work, Sub-SpaCE [25], which uses a genetic algorithm (GA) with custom mutation and initialization to generate counterfactual explanations by optimizing sparsity, plausibility and minimizing the number of subsequences. Multi-SpaCE inherits all the good properties of Sub-SpaCE (such as its model-agnostic nature, its computational efficiency or the validity of all the generated CFEs), while overcoming its two major limitations: its restriction to univariate data and its reliance on extensive experimentation to balance multiple desiderata. By introducing a multi-objective optimization framework based on Non-Dominated Sorting Genetic Algorithm II (NSGA-II) [26], Multi-SpaCE offers a Pareto-Front of solutions, enabling users to explore trade-offs between objectives. The optimization task identifies the points in the original input to modify, substituting their values with those of the Nearest Unlike Neighbor (NUN) [17]. Figure 1 presents the architecture of the proposed solution: the NUNs obtained for each input instance, the black box classifier to be explained, the NSGA algorithm used to generate candidate CFEs, the Autoencoder introduced to measure their plausibility, and the Pareto-front of counterfactuals returned by Multi-SpaCE.
Figure 1: Multi-SpaCE block diagram, including the black-box classifier to be explained, and the autoencoder used to measure the plausibility of the generated CFEs.
@@
To the best of our knowledge, Multi-SpaCE is the first multi-objective optimization method that ensures perfectly valid counterfactual explanations for multivariate time series classification problems. Extensive experimentation demonstrates its ability to achieve perfect validity across different datasets while outperforming existing methods in multiple metrics.
@@ 2 Related work
The field of eXplainable Artificial Intelligence (XAI) has experienced significant growth in recent years, resulting in the development of numerous methods tailored to specific domains and data types [27]. However, XAI methods for time series have received considerably less attention than those designed for tabular or image data [28, 29]. Many existing methods are adaptations of those designed for tabular or image data [30], and cannot properly cope with the unique characteristics of time-ordered inputs [31]. Time series data, such as electroencephalograms (EEG), electrocardiograms (ECG), human motion, or financial signals, require domain expertise to interpret, making traditional methods less effective to generate meaningful explanations [28, 32].
@@
This tendency extends to counterfactual explanations. Formally, a counterfactual explanation $x^{\prime}$ is the smallest variation of an original instance $x$ that changes the predicted outcome $y=b(x)$ of a black-box classifier $b\!:X\rightarrow[0,1]$, subject to $b(x^{\prime})\neq y$ [7]. Most counterfactual explanation methods are designed for tabular data. When applied to time series, these methods may not scale well with the increased dimensionality of inputs and often fail to account for the serial correlations that characterize time series data [17].
@@
To address these challenges, methods such as Native Guide (NG) [17] introduced contiguity, ensuring that changes occur as subsequences aligned with the sequential structure of time series. NG finds the Nearest Unlike Neighbor (NUN), the closest example to the instance being explained that belongs to a different class, and substitutes the shortest subsequence of the original instance with the corresponding subsequence from the NUN to achieve the desired outcome. The main drawback of NG is that it is limited to univariate settings and allows only a single subsequence of changes, thus restricting its practical applicability. COMTE, by Ates et al. [18], extended this idea to multivariate time series by optimizing the number of channels substituted from the NUN to produce a valid counterfactual. However, this approach is still limited to a single subsequence per channel, which may introduce unnecessary changes and break the existing relationships between channels.
@@
Other methods leverage shapelets to guide counterfactual explanations. For example, Local Agnostic Subsequence-based Time Series Explainer (LASTS) [19, 20] uses the latent space of a Variational Autoencoder (VAE) to generate a set of exemplars and counterfactuals in the neighborhood of the instance to explain. It then derives classification rules based on the presence of specific shapelets. Another work, Shapelet Explainer for Time Series (SETS) [33], extracts discriminative shapelets for each class. It then localizes them in the input to explain, and substitutes their corresponding values with those of the NUN.
@@
More recent approaches, such as Attention-Based Counterfactual Explanations (AB-CF) [22] and Discord-based Counterfactual Explanations (DiscoX) [23], address multivariate time series but rely on fixed-length subsequences. AB-CF greedily selects fixed-length multichannel subsequences that maximize classification entropy when isolated, while DiscoX identifies fixed-length discords (subsequences maximally distant from their nearest neighbors in the desired class) and replaces them with the nearest subsequences from the desired output class. Although these methods perform well in some scenarios, their reliance on fixed-length subsequences often results in unnecessary changes. Additionally, their heuristic and greedy approaches reduce their validity rates, limiting their applicability in real-world scenarios.
@@
Another line of research still follows the optimization-based counterfactual generation originally proposed by Wachter et al. [4]. Building on this foundational work, Glacier [21] generates counterfactuals in the latent space of an autoencoder while penalizing changes outside relevant subsequences identified by LIMESegment [34]. However, Glacier treats validity as a loss term rather than a strict constraint, leading to low validity rates, and remains limited to univariate datasets. Genetic algorithms (GAs) have also been applied to generate CFEs. TSEvo [16] introduced a multi-objective approach and diverse mutation operators to optimize both the location and magnitude of changes. However, TSEvo’s high-dimensional search space results in long execution times and low validity rates. More recently, TX-Gen [24] has tackled multi-objective optimization by finding a single subsequence of changes. Instead of relying on the NUN, TX-Gen employs a compressed representation of the subsequence and uses a simple AR model to generate the content of the subsequences proposed by the optimizer. Although TX-Gen reports perfect validity, it was only tested in univariate settings and might struggle with plausibility in more complex scenarios, where a simple generative AR model might not provide enough flexibility. Finally, Sub-SpaCE [25] balances sparsity, plausibility, and the number of subsequences by framing counterfactual generation as a single-objective optimization problem solved using a GA with custom mutation and initialization. Sub-SpaCE identifies a binary mask of changes, indicating the locations where the values of the original instance should be substituted with those of the NUN. While effective, Sub-SpaCE is restricted to univariate data and lacks a multi-objective framework, requiring experimentation to balance objective relevance.
@@
As stated above, contiguity is a desired property, as grouped changes in the form of subsequences enhance interpretability. However, current methods either promote contiguity indirectly (e.g., TSEvo by using the mutation operators), or limit their flexibility by imposing: i) a single subsequence of changes (e.g., NG, COMTE and TX-Gen) or; ii) fixed-length subsequences (e.g., AB-CF and DiscoX), thus hindering their adaptability to complex scenarios. Additionally, methods should be prepared to work in multivariate settings, which are ubiquitous in real-world applications, something that NG, Glacier, TX-Gen, and Sub-SpaCE do not support. Finally, and most importantly, except for a few approaches (such as NG, TX-Gen and Sub-SpaCE), most methods fail to ensure perfect validity, treating it as another objective rather than as a strict constraint.
@@ 3.2.1 Extension to multivariate problems
In both cases the NUN is the same as in Sub-SpaCE: the instance, $\mathbf{x}^{nun}$, closest to $\mathbf{x}$ in Euclidean space. Both approaches have been tested, and the best results (see Appendix A) are obtained when both are used: the common mask across channels is used at the beggining of the optimization, and the independent mask is used to then improve the solution. The pseudocode of Multi-SpaCE is shown in Algorithm 1.
@@ 4.1 Setup
For each dataset in Table 1, we trained an InceptionTime [35], a state-of-the-art model to perform classification with time series data. Additionally, we trained several Autoencoders ($f_{AE}$) to assess the plausibility of the generated counterfactuals, as well as two alternative models: Isolation Forests ($f_{IF}$), and Local Outlier Factor ($f_{LOF}$) as used in [23]. Details about the trained models are provided in the Supplementary Material at https://github.com/MarioRefoyo/Multi-SpaCE. We then selected the best-performing models of each family. For the Autoencoder, the one with the lowest reconstruction error was chosen. Models with the best silhouette scores were used for the Isolation Forest and Local Outlier Factor. Classifiers and autoencoders were implemented using the TensorFlow framework, while Isolation Forests and Local Outlier Factors were trained using the Scikit-Learn library.
The parameter values of Multi-SpaCE were empirically set and kept consistent across datasets to ensure robustness in multiple scenarios. Instead of extensive fine-tuning of all possible parameters, we focused on evaluating the influence of mutation parameters and the different types of masks proposed in Section 3.2.1, as they are central to our framework. Based on these evaluations (see A), the final Multi-SpaCE uses a population size $N=100$ and a total number of generations $G=100$, with the first $G_{1}=75$ generations used to optimize a common mask across channels with $p^{e}=p^{c}=0.75$ and $p^{p}=0$, and the last $G_{2}=25$ generations used to prune unnecessary subsequences with the independent mask setting with $p^{p}=0.75$. We use an initialization percentage $h=20\%$, increased by $h_{inc}=20\%$ if no valid counterfactual is found after $G_{reinit}=50$ generations. The penalization term for incorrect classes was set to $\nu=100$ to strictly enforce valid counterfactuals. The implementation of Multi-SpaCE and experimental code is publicly available at https://github.com/MarioRefoyo/Multi-SpaCE.
@@
Unlike the baseline methods, which return a single counterfactual considered optimal according to their specific criteria, Multi-SpaCE generates a set of solutions that are optimal under different weightings of the objectives in (5). To compare Multi-SpaCE with the baselines, we use the utility function proposed in Sub-SpaCE [25] to assign weights to each objective: a weight of 0.1 to the adversarial objective ($o_{1}$), 0.3 to sparsity ($o_{2}$), 0.4 to the number of subsequences ($o_{3}$), and 0.2 to plausibility ($o_{4}$). After applying the utility function, we select the counterfactual from the Multi-SpaCE solution set that maximizes the utility score. This selected counterfactual is then used to compute the metrics, with the results reported in Section 4.5.
@@ 4.2 Datasets
The performance of Multi-SpaCE was evaluated on datasets from the UCR archive [36] and the UEA archive [37]. Only datasets where the classifier achieved an F1-score greater than 80% were considered, ensuring that counterfactuals were generated for reasonably accurate models.
Under the multivariate setting, only 10 datasets surpassed the 80% threshold. For the univariate setting, over 40 datasets met the criterion, out of which 15 were selected to ensure diversity based on the following criteria: i) number of classes; ii) time series length; and iii) different data characteristics. In this regard, half of the chosen datasets had a binary classification outcome, as some baseline methods are tailored for binary classification. Additionally, we selected extreme cases, such as NonInvasiveFatalECGThorax2, which has a large number of classes (42 classes). Datasets with varying lengths were included, such as ItalyPowerDemand (24 steps) and HandOutlines (2709 steps). Finally, we selected datasets with diverse characteristics. For instance, FordA was chosen since frequencies are representative of output classes, while datasets like ECG200 and TwoPatterns were included as waveform patterns are what define class separability. Table 1 provides a detailed listing of all datasets (train and test size, time series length, number of channels and classes), as well as the test F1-scores of the trained classifiers.
@@
Coffee 28 28 1 286 2 1.000
Gunpoint 50 150 1 150 2 1.000
Plane 105 105 1 144 7 1.000
TwoPatterns 1000 4000 1 128 4 1.000
CBF 30 900 1 128 3 0.999
Strawberry 613 370 1 235 2 0.981
FordA 3601 1320 1 500 2 0.959
ItalyPowerDemand (ItalyPower) 67 1029 1 24 2 0.958
FacesUCR 200 2050 1 131 14 0.951
HandOutlines 1000 370 1 2709 2 0.948
NonInvasiveFatalECGThorax2 (NI-ECG2) 1800 1965 1 750 42 0.948
ProximalPhalanxOutlineCorrect (PPOC) 600 291 1 80 2 0.913
ECG5000 500 4500 1 140 5 0.901
ECG200 100 100 1 96 2 0.876
CinCECGTorso 40 1380 1 1639 4 0.856
BasicMotions 40 40 6 100 4 1.000
PenDigits 7494 3498 2 8 10 0.988
Epilepsy 137 138 3 206 4 0.971
NATOPS 180 180 24 51 6 0.933
UWaveGestureLibrary (UWave) 120 320 3 315 8 0.883
RacketSports 151 152 6 30 4 0.875
ArticularyWordRecognition (AWR) 275 300 9 144 25 0.872
Cricket 108 72 6 1197 12 0.870
SelfRegulationSCP1 (SR-SCP1) 268 293 6 896 2 0.837
PEMS-SF 267 173 963 144 7 0.813